The length of a wire bent intoa rectangle is twice its breadth. If this rectangle is turned into a circle, the circumference becomes 77 cm. Find the length and breadth of the rectangle
Found 2 solutions by ilana, kirubakaran:
Answer by ilana(307)

The circumference of the circle is the same amount of wire as the perimeter of the rectangle. So we can say 2l+2w=77. We know l=2w, so we can substitute this into the first equation to get 2(2w)+2w=77, So 4w+2w=77, so 6w=77. So the width (or breadth) is 77/2cm, and the length, twice its breadth, is 77cm.

Answer by kirubakaran(5)

The length of a wire bent intoa rectangle is twice its breadth. If this rectangle is turned into a circle, the circumference becomes 77 cm. Find the length and breadth of the rectangle
circumference of a circle = 2 x 22/7 x r = 77cm
r= 77 x 7/22 x 1/2
r= 49/4
2l + 2b= 2x22/7x r
= 2x 22/7 x 49/4
2l + 2b= 77
2(2b)+ 2b = 77 Length is twice its breadth. 4b + 2b = 77
b = 77/6
l = 2b = 77/3